### v1

Abstract
Quantum machine learning offers a promising pathway for enhancing stock market prediction, particularly under complex, noisy, and highly dynamic financial environments. However, many classical forecasting models struggle with noisy input, regime shifts, and limited generalization capacity. To address these challenges, we propose a Quantum Temporal Convolutional Neural Network (QTCNN) that combines a classical temporal encoder with parameter-efficient quantum convolution circuits for cross-sectional equity return prediction. The temporal encoder extracts multi-scale patterns from sequential technical indicators, while the quantum processing leverages superposition and entanglement to enhance feature representation and suppress overfitting. We conduct a comprehensive benchmarking study on the JPX Tokyo Stock Exchange dataset and evaluate predictions through long-short portfolio construction using out-of-sample Sharpe ratio as the primary performance metric. QTCNN achieves a Sharpe ratio of 0.538, outperforming the best classical baseline by approximately 72%. These results highlight the practical potential of quantum-enhanced forecasting model, QTCNN, for robust decision-making in quantitative finance.
Index Terms:
Quantum Machine Learning, Financial Time-Series Forecasting, Temporal Convolutional Neural Network
[…]
=-15pt
I Introduction
Stock market forecasting is a major challenge in computational finance[10545224]. Financial markets produce huge amounts of complex data every day. This data is noisy, changes constantly, and contains many hidden patterns[10862401]. Creating models that accurately predict market movements is highly valuable, as strong forecasts enable better investment strategies and more effective risk management.
In recent years, researchers have used many different computational methods to predict stock prices. These include classic statistical models [11115662], machine learning algorithms, and modern deep learning methods like Transformers [9357288]. While these methods have shown some success, they often face significant problems. Many models struggle to capture the complex, long-term dynamics of the market. They may also overfit to past data, which means they perform poorly when market conditions change[11140787]. As a result, many existing methods lack the robustness and adaptability needed for real-world trading.
Quantum computing offers a new and promising approach to these challenges[10461108]. Quantum Machine Learning (QML) uses the principles of quantum mechanics, such as superposition and entanglement. These features allow quantum models to explore very large, complex data spaces[10864082]. This may help them find important patterns in financial data that classical computers cannot.
Recent studies have explored a diverse range of computational and intelligent approaches to improve the accuracy and efficiency of stock market forecasting. These works span from deep learning architectures and hybrid statistical-machine learning frameworks to quantum-inspired algorithms.
Aşırım, et al.[11171624] employed a single-layer Transformer model to forecast daily closing prices of Nasdaq stocks. By segmenting and independently normalizing historical data with sinusoidal positional embeddings, the model achieved an average correlation of about 0.96, effectively and efficiently capturing stock price trends.
Ahire et al.[10882429] proposed a hybrid Auto Regressive Integrated Moving Average (ARIMA) and machine learning model to forecast stock trends using ICICI and SBI data. By combining ARIMA residuals with technical indicators such as Relative Strength Index (RSI), Simple Moving Average (SMA), and Stochastic K, their method improved prediction accuracy to about 0.96 for SBI and 0.94 for ICICI, effectively identifying bullish and bearish movements.
Singh et al.[10545132] applied multiple machine learning algorithms to predict stock performance of HRM firms Info Edge Ltd and Quess Corporation using five years of data. Results showed that Random Forest and Gradient Boosting achieved the highest accuracy, with $R^{2}$ values above 0.999 and Mean Absolute Percentage Error (MAPE) below 0.2%.
RaviTeja et al.[10690751] developed a Novel Quantum Enforcing Algorithm (NQEA) to forecast stock index and Bitcoin prices, comparing it with the Bernstein-Vazirani Algorithm (BV). Using datasets with 1,400 training and 1,856 testing samples. Results showed that NQEA achieved an accuracy of 95.9% with a loss of 4.1%, outperforming BV’s 88.0% accuracy and 12.0% loss.
Hossain et al.[11171862] applied Linear Regression, Random Forest, Gradient Boosting, XGBoost, and Multilayer Perceptron (MLP), to predict stock prices of Dhaka Stock Exchange. Results showed that Linear Regression achieved the lowest prediction error and most consistent performance across all datasets, outperforming ensemble and neural network models.
Despite their notable results, the studies above share several common weaknesses. Most relied on limited dataset quality, which restricted generalization across diverse market conditions. The Transformer and several traditional machine learning approaches still struggled to capture nonlinear and long-term dependencies within financial time-series data, often showing risks of overfitting and limited incorporation of richer signals such as macroeconomic or sentiment-driven factors.
In this paper, we propose a quantum-enhanced approach to address these limitations. Our model QTCNN, which leverages superposition and entanglement to extract high-dimensional, nonlinear relationships within noisy market data. This architecture is designed to improve generalization and boost robustness compared with classical models by providing richer representational capacity while maintaining computational efficiency.
To validate our model, we conduct a comprehensive benchmarking study using the JPX Tokyo Stock Exchange dataset[KaggleJPX2025]. We compare our proposed QTCNN model against a strong set of classical baselines, including LightGBM (LGBM)[NIPS2017_6449f44a], LSTM[11064676], and the Transformer[10920805]. We also compare it to other recently published quantum-inspired algorithms. Our results show that our model achieves superior predictive accuracy and robustness, demonstrating a more effective way to handle complex financial time-series data.
The main contributions of this paper can be concluded as bellow:
  • We proposed an end-to-end quantum-enhanced framework for cross-sectional equity return prediction, where the QNN serves as the primary predictive model to extract complex market signals.
  • The QNN leverages quantum superposition and entanglement to model nonlinear, high-dimensional structures in noisy financial data, improving generalization capability while mitigating overfitting compared with classical machine learning approaches.
  • A unified benchmarking study is conducted against representative deep learning architectures, statistical methods, and other quantum-based models, demonstrating the superior predictive accuracy and robustness of the proposed framework in a realistic financial forecasting scenario.
The rest of this article is organized as follows. Section II reviews the task and dataset used in this work. Section III details the specific architecture of our proposed QTCNN model. Section IV describes the experimental setup and presents our numerical results. Finally, Section V concludes this paper and discusses potential directions for future research.
II Task and Dataset
II-A JPX Tokyo Stock Exchange Prediction
This section describes the official dataset of the Kaggle "JPX Tokyo Stock Exchange Prediction" competition and how it is used in our study, including the construction of training labels and subset sampling schemes for computationally efficient yet reproducible experiments.
II-B Task Overview
  • Prediction task. For each trading day, the model produces a cross-sectional ranking of stocks. Based on this ranking, we form a long–short portfolio by going long the top $N$ stocks and short the bottom $N$ stocks.
  • Model output. On each trading day, the model outputs a continuous score (or probability) for every stock. These scores are standardized cross-sectionally (within the day) and then used to rank the stocks.
  • Performance metric. We evaluate the model by computing the Sharpe ratio of the long–short portfolio over an out-of-sample (OOS) period, using the daily portfolio returns implied by the official competition target.
II-C Original Dataset and Stock Universe
II-C1 Data source and time span
The dataset is provided by JPX via the Kaggle competition. The official training data (“train” stock prices) covers approximately 2017–2021 Japanese trading days, and the supplemental files extend the timeline into 2022. On each trading day, there are roughly $\sim 2000$ “active” stocks with complete price, volume, and target information.
II-C2 Universe definition and Universe0
In addition to price data, JPX provides a static stock information table, denoted as stock_list, which contains a key flag Universe0:
  • Universe0 = 1 (or True) identifies stocks that belong to a core universe of high-capitalization, sufficiently liquid names (approximately the top $\sim 2000$ stocks by size and liquidity).
  • These Universe0 stocks constitute the official prediction universe in the competition.
In our study, we focus on stocks with Universe0 = True to match the official setting and to avoid extremely illiquid securities.
II-D Key Fields and Financial Interpretation
II-D1 Price data (stock_prices)
The main price table (both in the training and supplemental sets) contains the following fields used in this work:
  • Date: Trading date.
  • SecuritiesCode: Stock identifier (e.g., 1301, 1332), serving as a key together with Date.
  • Open, High, Low, Close: Daily open, high, low, and close prices.
  • Volume: Daily traded volume (in shares).
  • AdjustmentFactor: Adjustment factor capturing stock splits, reverse splits, share distributions, and other corporate actions.
  • SupervisionFlag: Indicator for stocks under special supervision by the exchange or at risk of delisting.
  • Target: The official competition target, corresponding to a two-day holding return defined on future prices (see Section II-E).
These fields form the raw basis for our feature engineering and financial interpretation.
II-D2 Static stock information (stock_list)
The stock_list table provides relatively static attributes for each stock:
  • SecuritiesCode: Stock identifier, matching stock_prices.
  • Name: Stock name.
  • MarketCapitalization: Market capitalization.
  • Universe0: Boolean flag indicating membership in the core prediction universe.
In the present work, we primarily use Universe0 to define the prediction universe. Sector codes are not used as input features in the main models but can be exploited for industry-neutral analysis or robustness checks.
II-D3 Adjusted close price
To eliminate non-economic jumps in price series caused by corporate actions, we construct an adjusted close price for each stock $k$ at time $t$ using the cumulative adjustment factors:
$\mathrm{AdjClose}_{k,t}=\mathrm{Close}_{k,t}\cdot\prod_{\tau\leq t}\mathrm{AdjustmentFactor}_{k,\tau}.$ (1)
Here, $k$ indexes the stock, and $t$ indexes the trading day. The adjusted close $\mathrm{AdjClose}_{k,t}$ is used to compute returns, momentum, volatility, and other time-series features.
Importantly, the official Target remains defined on the raw closes; we do not modify the target. Adjustments are applied only on the feature side.
II-E Definition and Interpretation of the Target
The Target field is the central label of the JPX competition and corresponds to a two-day holding return with a one-day execution delay.
Let $C(k,t)$ denote the raw close price of stock $k$ at trading day $t$. The official target can be written as:
$\mathrm{Target}(k,t)=\frac{C(k,t+2)-C(k,t+1)}{C(k,t+1)}.$ (2)
Financially, this corresponds to the following strategy:
  • On day $t$, the model observes information up to and including day $t$ and produces a ranking for day $t$.
  • At the close of day $t+1$, a position is opened at price $C(k,t+1)$ (long or short depending on the ranking).
  • At the close of day $t+2$, the position is closed at price $C(k,t+2)$.
  • The realized return is exactly $\mathrm{Target}(k,t)$ as defined above.
Thus, the model at day $t$ predicts the return of a two-day holding strategy executed over $[t+1,t+2]$. This design enables daily rebalancing while avoiding intraday trading assumptions.
II-F Derived Features and Preprocessing
II-F1 Single-stock time-series features
Starting from $\mathrm{AdjClose}$, we compute for each stock $k$ a set of univariate time-series features:
  • Returns and momentum
    • One-period return:
      $\mathrm{ret1}_{k,t}=\dfrac{\mathrm{AdjClose}_{k,t}-\mathrm{AdjClose}_{k,t-1}}{\mathrm{AdjClose}_{k,t-1}}.$ (3)
    • Momentum indicators $\mathrm{mom}\{1,2,5,10,20\}$, e.g., cumulative return over the past $n$ days or deviations from moving averages.
  • Volatility features
    • Rolling standard deviation of returns: $\mathrm{vol}\{5,10,20\}$ computed from $\mathrm{ret1}$ over windows of length $5$, $10$, and $20$ days.
  • Price–volume features
    • Dollar volume: $\mathrm{dv}_{k,t}=\mathrm{Close}_{k,t}\times\mathrm{Volume}_{k,t}.$
    • Rolling means of volume and dollar volume: $\mathrm{dv\_mean}\{5,10,20\}$, $\mathrm{vol\_mean}\{5,10,20\}$.
    • Log-transformed features such as
      $\log(\mathrm{Volume}_{k,t})$ and $\log(\mathrm{dv}_{k,t})$ to mitigate heavy-tailed distributions.
II-F2 Cross-sectional normalization
For each trading day $t$, we apply a cross-sectional preprocessing pipeline over all stocks (typically restricted to Universe0):
  1. 1.
    Winsorization. For each feature, values are winsorized at the 11st and $99$th percentiles across stocks on day $t$ to trim extreme outliers.
  2. 2.
    Z-score standardization. For winsorized feature values $\{x_{k,t}\}_{k}$, we compute
    $z_{k,t}=\frac{x_{k,t}-\mu_{t}}{\sigma_{t}},$ (4)
    where $\mu_{t}$ and $\sigma_{t}$ are the cross-sectional mean and standard deviation of that feature on day $t$.
This procedure emphasizes relative signals within each day’s cross-section rather than absolute price or volume levels, aligning with the cross-sectional stock selection framework.
II-G Label Construction and Weakly Supervised Training Samples
Although the official Target is a continuous value, we adopt a weakly supervised scheme that focuses on the most clearly positive and negative examples.
For each trading day $t$:
  1. 1.
    We rank all stocks by $\mathrm{Target}(k,t)$ in descending order.
  2. 2.
    We select:
    • The top $p$ stocks as the positive class with binary label $\mathrm{label\_bin}(k,t)=1$;
    • The bottom $p$ stocks as the negative class with binary label $\mathrm{label\_bin}(k,t)=0$.
  3. 3.
    Stocks in the middle range are excluded from training and treated as low-signal observations.
In our experiments, $p$ is typically set to $200$ or $300$. This construction increases the signal-to-noise ratio by concentrating supervision on the strongest long and short candidates and mirrors the eventual long–short portfolio structure used in evaluation.
II-H Temporal Split and Sequence Construction
II-H1 Temporal train–test split
To avoid look-ahead bias, we perform a simple chronological split:
  • After sorting all days by Date, the first $80\%$ of days constitute the training period.
  • The remaining $20\%$ of days constitute the out-of-sample test period.
All model training and hyperparameter tuning are conducted on the training period. The Sharpe ratios reported in our results are computed solely on the OOS period.
II-H2 Sequence construction for time-series models
For sequence models such as LSTMs, Transformers, or quantum-enhanced variants, we construct input sequences of fixed length $\mathrm{seq\_len}$ (default $\mathrm{seq\_len}=20$):
$\mathbf{x}_{k,t}=\big[\mathrm{features}_{k,t-19},\ldots,\mathrm{features}_{k,t}\big],$ (5)
where $\mathrm{features}_{k,\tau}$ denotes the feature vector for stock $k$ at day $\tau$. Each sequence $\mathbf{x}_{k,t}$ is paired with the label $\mathrm{Target}(k,t)$ or its binary counterpart $\mathrm{label\_bin}(k,t)$.
II-I Subset Sampling of Trading Days
To control computational cost and facilitate fair comparisons across different model architectures, we design two sampling schemes at the trading-day level. In both schemes, sampling is applied before feature engineering and sequence construction.
II-I1 Stride sampling
Let $\{t_{1},t_{2},\ldots,t_{T}\}$ be the chronologically ordered trading days. Given a stride $k$ (e.g., $k=11$), we select the subset
$\{t_{1},t_{1+k},t_{1+2k},\ldots\}.$ (6)
This “stride” subsampling:
  • Preserves the overall temporal span (e.g., 2017–2021) while sparsifying the calendar.
  • Uses approximately $1/k$ of all trading days (for $k=11$, about $9\%$ of the original days).
  • Substantially reduces data volume and training time, making it suitable for rapid prototyping and architectural exploration.
II-I2 Year-wise stratified fraction sampling
Alternatively, we partition trading days by calendar year (e.g., 2017, 2018, …, 2021) and perform stratified random sampling:
  1. 1.
    For each year, we randomly sample a fixed fraction $\alpha$ (e.g., $\alpha=0.1$) of trading days from that year.
  2. 2.
    A fixed random seed is used to ensure reproducibility.
This year-wise stratification:
  • Ensures that different market regimes (bull, sideways, crisis periods, etc.) are proportionally represented in the subset.
  • Avoids concentrating the subset on a single year with exceptionally high volatility or data density.
II-I3 Relation between subsets and the full dataset
All subsequent steps—feature engineering, label construction, temporal splitting, and sequence construction—are performed on the selected subset of trading days. Using the full dataset is equivalent to setting the stride to $k=1$ or the fraction to $\alpha=1$ (i.e., no subsampling).
These elements fully specify how the original JPX dataset is transformed into the training and evaluation sets used in our experiments.
III Methodology
III-A Quantum State Representation
An $n$-qubit quantum state resides in the Hilbert space $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$ with dimension $2^{n}$. A pure state is represented as:
[…]
III-B Parameterized Quantum Circuits
Variational quantum algorithms employ parameterized unitary operations $U(\bm{\theta})$ where $\bm{\theta}\in\mathbb{R}^{p}$ are trainable parameters [cerezo2021variational]:
$U(\bm{\theta})=\prod_{l=1}^{L}W_{l}U_{l}(\bm{\theta}_{l})$ (8)
where $W_{l}$ are fixed entangling operations and $U_{l}(\bm{\theta}_{l})$ are parameterized rotations.
III-C Feature Encoding
Classical data $\bm{x}\in\mathbb{R}^{n}$ is encoded into quantum states via angle embedding [schuld2021machine]:
[…]
where $R_{Y}(\theta)=\exp(-i\theta Y/2)$ is the Pauli-Y rotation gate.
III-D Measurement and Expectation Values
Predictions are obtained by measuring Pauli observables:
[…]
For binary classification, the Pauli-Z expectation $\langle Z\rangle\in[-1,1]$ on designated qubits provides the model output.
III-E Quantum Temporal Convolutional Neural Network (QTCNN)
The central methodological contribution of this work is the Quantum Temporal Convolutional Neural Network (QTCNN), which addresses two key limitations of existing quantum approaches for sequential data:
  1. 1.
    Temporal feature extraction: Unlike standard QCNN that applies PCA to flatten temporal structure, QTCNN employs a classical temporal encoder to preserve sequential dynamics before quantum processing.
  2. 2.
    Parameter efficiency: Through intra-layer parameter sharing, QTCNN reduces quantum circuit parameters from $O(n_{q}\cdot L)$ to $O(L)$, mitigating trainability issues associated with barren plateaus in deep variational circuits.
We introduce the QTCNN, a hybrid quantum-classical architecture specifically tailored for financial time series data. The QTCNN architecture implements quantum convolutional operations inspired by classical CNNs [cong2019quantum, henderson2020quanvolutional].
The remaining quantum architectures (QNN, QCNN, QLSTM, QASA, QSVM, PQC+LGBM) serve as ablation baselines to isolate the contribution of each design choice.
III-E1 Input Processing
Tabular features $\bm{x}\in\mathbb{R}^{d}$ undergo dimensionality reduction:
$\bm{z}=W_{\text{PCA}}\cdot\text{StandardScaler}(\bm{x})\in\mathbb{R}^{n_{q}}$ (11)
where $n_{q}=8$ qubits and $W_{\text{PCA}}\in\mathbb{R}^{n_{q}\times d}$ is the PCA projection matrix.
III-E2 Quantum Convolutional Layer
The QTCNN circuit applies hierarchical convolution-pooling operations. The convolutional unitary on adjacent qubits $(i,i+1)$ is:
$U_{\text{conv}}(\bm{\theta})=\text{CNOT}_{i,i+1}\cdot(R_{Y}(\theta_{1})\otimes R_{Z}(\theta_{2}))\cdot$ (12)
$\text{CNOT}_{i+1,i}\cdot(R_{Y}(\theta_{3})\otimes R_{Z}(\theta_{4}))\cdot(R_{Y}(\theta_{5})\otimes R_{Z}(\theta_{6}))$
with 6 parameters per convolutional unit. After each convolutional layer, pooling is performed by measuring and discarding alternating qubits, reducing the circuit width by half.
III-E3 Hybrid Classification Head
The quantum output $q\in\mathbb{R}$ (single-qubit measurement) is concatenated with classical features:
$\hat{y}=\sigma\left(\text{MLP}([q;\bm{z}])\right)$ (13)
where $\text{MLP}:\mathbb{R}^{n_{q}+1}\to\mathbb{R}$ has architecture $64\to 32\to 1$ with ReLU activations, and $\sigma(\cdot)$ is the sigmoid function.
III-E4 Training Objective
The model minimizes binary cross-entropy loss:
$\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right]$ (14)
III-F Quantum Temporal Convolutional Neural Network (QTCNN) vs. QCNN
Overview
QTCNN adopts a hybrid architecture comprising a classical temporal encoder, a simplified QCNN, and a classical MLP. In contrast to conventional QCNN approaches that directly apply angle embedding to dimensionality-reduced vectors, QTCNN first employs a temporal CNN to extract an $n_{q}$-dimensional representation from a $T\times F$ feature sequence, which is subsequently fed into a hierarchical quantum convolutional and pooling structure with parameter sharing. The final classification scores are obtained by concatenating the quantum output with classical feature vectors.
III-F1 Input Processing
QCNN (baseline). Tabular features $\bm{x}\in\mathbb{R}^{d}$ are standardized and projected to $n_{q}$ dimensions via PCA:
$\bm{z}_{\text{QCNN}}$ $=\bm{W}_{\text{PCA}}\cdot\text{StandardScaler}(\bm{x}),$ (15)
$\bm{z}_{\text{QCNN}}$ $\in\mathbb{R}^{n_{q}},\quad\bm{W}_{\text{PCA}}\in\mathbb{R}^{n_{q}\times d}.$
QTCNN (ours). A temporal sample $X\in\mathbb{R}^{T\times F}$ composed of multi-scale technical factors is compressed to $n_{q}$ dimensions through a temporal CNN encoder $f_{\text{TConv}}$:
$\bm{z}_{\text{QTCNN}}=f_{\text{TConv}}(X;\,\bm{\phi})\in\mathbb{R}^{n_{q}},$ (16)
where $f_{\text{TConv}}$ consists of 1D convolutions followed by global average pooling (GAP) over the temporal dimension and an MLP, thereby preserving sequential dynamics.
III-F2 Quantum Embedding
Both approaches employ angle embedding to encode features into the quantum circuit; this work utilizes $R_{Y}$ rotations:
$\text{AngleEmbedding}(\bm{z},\text{rotation}=Y),\quad\bm{z}\in\mathbb{R}^{n_{q}}.$ (17)
III-F3 Quantum Convolution and Pooling
The quantum convolutional unit operates on adjacent qubit pairs (i,i+1)(i,i{+}1) as follows:
$U_{\text{conv}}(\bm{\theta})$ $=\text{CNOT}_{i,i+1}\cdot(R_{Y}(\theta_{1})\otimes R_{Z}(\theta_{2}))\cdot\text{CNOT}_{i+1,i}$ (18)
$\quad\cdot(R_{Y}(\theta_{3})\otimes R_{Z}(\theta_{4}))\cdot(R_{Y}(\theta_{5})\otimes R_{Z}(\theta_{6})),$
where each convolutional unit contains $6$ trainable parameters.
Key Distinction: Parameterization and Depth Control
  • QCNN (baseline). The conventional approach assigns independent parameters $\bm{\theta}^{(\ell)}_{i}$ to each adjacent pair at layer $\ell$.
  • QTCNN (ours). We adopt intra-layer parameter sharing: a single parameter set $\bm{\theta}^{(\ell)}\in\mathbb{R}^{6}$ at layer $\ell$ is shared across all adjacent pairs within that layer, substantially reducing the parameter count and improving training stability.
Pooling halves the circuit width through interleaved discarding (e.g., retaining even-indexed qubits):
$\mathcal{W}_{\ell+1}=\text{Pool}(\mathcal{W}_{\ell})=\mathcal{W}_{\ell}[::2],$ (19)
with the effective depth automatically constrained to $L_{\text{eff}}=\min\!\big(L,\lfloor\log_{2}(n_{q})\rfloor\big)$ to prevent excessive pooling that would leave no measurable qubits.
III-F4 Hybrid Classification Head
A single retained qubit is measured to yield the quantum output $q\in\mathbb{R}$, which is then concatenated with the classical feature vector to produce the final output:
$\hat{y}=\sigma\!\left(\text{MLP}\big([\,q;\,\bm{z}\,]\big)\right),\quad$ (20)
$\text{MLP}:\mathbb{R}^{n_{q}+1}\to\mathbb{R},\ \ 64\to 32\to 1\ \text{with ReLU}.$ (21)
In QTCNN, $\bm{z}$ originates from the temporal encoder; in QCNN, $\bm{z}$ is derived from PCA.
III-F5 Training Objective
Both models are trained using binary cross-entropy loss:
$\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right].$ (22)
In practice, this work employs cross-sectional “extreme samples”—the top and bottom $K$ assets per day based on the target variable—as supervision signals to better approximate ranking and stock selection scenarios.
Summary of Differences
  • Feature pathway: QCNN reduces dimensionality from tabular features via PCA; QTCNN extracts a semantically richer representation $\bm{z}$ from $T\times F$ sequences using a temporal CNN.
  • Quantum convolution parameterization: QCNN typically employs pair-wise independent parameters; QTCNN adopts intra-layer parameter sharing with only $6$ parameters per layer, yielding substantial parameter reduction.
  • Depth and stability: QTCNN enforces $L_{\text{eff}}\leq\lfloor\log_{2}(n_{q})\rfloor$ to prevent over-pooling; the implementation processes samples individually in the forward pass to enhance numerical stability.
  • Fusion strategy: Both approaches concatenate $q;\bm{z}]$ and pass the result through an MLP; however, the $\bm{z}$ in QTCNN carries temporal semantics, which typically yields superior performance in ranking-based tasks.
III-G Quantum Long Short-Term Memory (QLSTM)
The QLSTM replaces classical gate computations with quantum circuits [chen2022quantum, di2022dawn].
III-G1 Sequence Embedding
Input sequences $\bm{X}=[\bm{x}_{1},\ldots,\bm{x}_{T}]\in\mathbb{R}^{T\times F}$ with $T=20$ are embedded:
$\bm{e}_{t}=\tanh\left(\text{LayerNorm}(W_{e}\bm{x}_{t}+b_{e})\right)\in\mathbb{R}^{h}$ (23)
where $h=64$ is the hidden dimension.
III-G2 Quantum Gate Computation
At timestep $t$, the concatenated vector $\bm{e}_{t};\bm{h}_{t-1}]\in\mathbb{R}^{2h}$ is processed by a quantum layer:
$\bm{u}_{t}$ $=\text{LayerNorm}(W_{q}[\bm{e}_{t};\bm{h}_{t-1}])\in\mathbb{R}^{n_{q}}$ (24)
$\bm{v}_{t}$ $=\text{QLayer}(\bm{u}_{t};\bm{\theta}_{q})\in\mathbb{R}^{n_{q}}$ (25)
where QLayer implements angle embedding followed by $L=2$ BasicEntanglerLayers:
$\text{QLayer}(\bm{u};\bm{\theta})=\left[\langle Z_{1}\rangle,\ldots,\langle Z_{n_{q}}\rangle\right]^{T}$ (26)
with circuit:
[…]
III-G3 LSTM Update Equations
The quantum output is projected to gate values:
$\bm{i}_{t};\bm{f}_{t};\bm{g}_{t};\bm{o}_{t}]=W_{g}\bm{v}_{t}+b_{g}\in\mathbb{R}^{4h}$ (28)
Standard LSTM updates follow:
$\bm{c}_{t}$ $=\sigma(\bm{f}_{t})\odot\bm{c}_{t-1}+\sigma(\bm{i}_{t})\odot\tanh(\bm{g}_{t})$ (29)
$\bm{h}_{t}$ $=\sigma(\bm{o}_{t})\odot\tanh(\bm{c}_{t})$ (30)
The final hidden state $\bm{h}_{T}$ is mapped to a logit via $W_{\text{out}}\bm{h}_{T}$.
III-H Variational Quantum Neural Network (QNN)
The QNN implements a pure quantum classifier with minimal classical post-processing [schuld2020circuit].
III-H1 Feature Mapping
Features are scaled to $0,\pi]$ for optimal expressibility:
$\phi_{j}=\pi\cdot\text{MinMaxScaler}(z_{j})\in[0,\pi]$ (31)
III-H2 Variational Circuit
The ansatz combines angle embedding with hardware-efficient entanglement:
$U(\bm{\phi},\bm{\theta})=\prod_{l=1}^{L}\left[\text{CNOT}_{\text{ring}}\cdot\bigotimes_{j=1}^{n_{q}}R_{Y}(\theta_{l,j})\right]\cdot\bigotimes_{j=1}^{n_{q}}R_{Y}(\phi_{j})$ (32)
where $L=2$ layers and $\text{CNOT}_{\text{ring}}$ applies entanglement in a circular topology.
III-H3 Output Layer
All Pauli-Z expectations are measured and linearly combined:
$\hat{y}=\sigma\left(\bm{w}^{T}\langle\bm{Z}\rangle+b\right),\quad\langle\bm{Z}\rangle=[\langle Z_{1}\rangle,\ldots,\langle Z_{n_{q}}\rangle]^{T}$ (33)
III-I Quantum Kernel Support Vector Machine (QSVM)
The QSVM leverages quantum feature maps to construct kernel functions [schuld2019quantum, havlivcek2019supervised].
III-I1 Quantum Feature Map
Classical features are encoded into quantum states:
[…]
using the angle embedding $S(\bm{x})$ from Equation (9).
III-I2 Quantum Kernel
The kernel function measures state overlap via the fidelity:
[…]
III-I3 SVM Optimization
The dual optimization problem with precomputed kernel $K$ is:
$\max_{\bm{\alpha}}\sum_{i=1}^{N}\alpha_{i}-\frac{1}{2}\sum_{i,j=1}^{N}\alpha_{i}\alpha_{j}y_{i}y_{j}K(\bm{x}_{i},\bm{x}_{j})$ (36)
subject to $0\leq\alpha_{i}\leq C$ and $\sum_{i}\alpha_{i}y_{i}=0$, solved using sklearn’s SVC with balanced class weights.
III-J Quantum-Adaptive Self-Attention (QASA)
QASA integrates quantum processing into Transformer architectures [chen2025quantum, chen2025qasa, chen2025qtws].
III-J1 Transformer Encoder
Sequential inputs $\bm{X}\in\mathbb{R}^{T\times F}$ are processed through embedding and self-attention [vaswani2017attention, li2023quantum]:
$\bm{E}$ $=\tanh(\text{LayerNorm}(\bm{X}W_{e}))\in\mathbb{R}^{T\times d}$ (37)
$\bm{H}$ $=\text{TransformerEncoder}(\bm{E})\in\mathbb{R}^{T\times d}$ (38)
where $d=64$ and the encoder uses multi-head attention with $n_{\text{heads}}=4$.
III-J2 Quantum Enhancement Layer
At each timestep $t$, the representation is quantum-processed:
$\bm{u}_{t}$ $=\text{LayerNorm}(W_{\text{proj}}\bm{h}_{t})\in\mathbb{R}^{n_{q}}$ (39)
$\bm{q}_{t}$ $=\text{QLayer}(\bm{u}_{t};\bm{\theta}_{q})\in\mathbb{R}^{n_{q}}$ (40)
$\tilde{\bm{h}}_{t}$ $=W_{\text{back}}\bm{q}_{t}\in\mathbb{R}^{d}$ (41)
The quantum layer applies the same ansatz as in QLSTM (Equation 26).
III-J3 Classification Output
The final timestep representation produces the prediction:
$\hat{y}=\sigma(W_{\text{out}}\tilde{\bm{h}}_{T}+b_{\text{out}})$ (42)
III-K PQC-Enhanced Gradient Boosting (PQC+LGBM)
This two-stage architecture combines quantum feature extraction with gradient boosting [chen2016xgboost, ke2017lightgbm].
III-K1 Stage 1: PQC Embedding
A parameterized quantum circuit is trained for regression:
$\hat{r}=W_{\text{reg}}^{T}\langle\bm{Z}\rangle_{\text{PQC}}+b_{\text{reg}}$ (43)
minimizing Huber loss:
$\mathcal{L}_{\text{Huber}}(r,\hat{r})=\begin{cases}\frac{1}{2}(r-\hat{r})^{2}&\text{if }|r-\hat{r}|\leq\delta\\ \delta(|r-\hat{r}|-\frac{\delta}{2})&\text{otherwise}\end{cases}$ (44)
III-K2 Stage 2: LambdaRank Optimization
Quantum embeddings $\langle\bm{Z}\rangle_{\text{PQC}}$ are used as features for LightGBM Ranker with LambdaRank objective [burges2010ranknet]:
$\mathcal{L}_{\text{LambdaRank}}=\sum_{(i,j):y_{i}>y_{j}}|\Delta\text{NDCG}_{ij}|\log(1+e^{-\sigma(s_{i}-s_{j})})$ (45)
where $s_{i},s_{j}$ are model scores and $\Delta\text{NDCG}_{ij}$ measures ranking quality change from swapping items $i$ and $j$.
III-L Classical Baseline Models
III-L1 Multi-Layer Perceptron (MLP)
A deep feedforward network:
$\hat{y}=\sigma\left(W_{3}\cdot\text{BN}\left(\text{ReLU}(W_{2}\cdot\text{BN}(\text{ReLU}(W_{1}\bm{x})))\right)\right)$ (46)
with hidden dimensions $[384,192,96]$, BatchNorm (BN), and Dropout ($p=0.1$).
III-L2 Long Short-Term Memory (LSTM)
Standard LSTM [hochreiter1997long] with embedding layer:
$\bm{h}_{T}=\text{LSTM}(\text{Embed}(\bm{X}))_{T},\quad\hat{y}=\sigma(W_{\text{out}}\bm{h}_{T})$ (47)
III-L3 Transformer Encoder
Following [vaswani2017attention]:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (48)
with two encoder layers, $d_{\text{model}}=64$, $n_{\text{heads}}=4$, and sinusoidal positional encoding.
III-L4 LightGBM Ranker
Gradient boosted decision trees with LambdaRank objective (Equation 45), configured with 1200 estimators, 63 leaves, and $L_{2}$ regularization $\lambda=2.0$.
III-L5 Momentum-Volatility Baseline
A parameter-free baseline:
$s^{(i)}=\frac{0.5\cdot\text{mom}_{5}^{(i)}+0.5\cdot\text{mom}_{20}^{(i)}}{\text{vol}_{20}^{(i)}+\epsilon}$ (49)
implementing risk-adjusted momentum [daniel2016momentum].
III-M Training Configuration
III-M1 Optimization Details
All neural network models are trained using gradient-based optimization. For quantum-classical hybrid models, parameter-shift rules [mitarai2018quantum, schuld2019evaluating] enable exact gradient computation:
$\frac{\partial\langle O\rangle}{\partial\theta_{j}}=\frac{1}{2}\left[\langle O\rangle_{\theta_{j}+\pi/2}-\langle O\rangle_{\theta_{j}-\pi/2}\right]$ (50)
III-M2 Hyperparameter Summary
TABLE I: Training Configuration Summary
Model Epochs Batch LR Optimizer Loss
QTCNN 50 128 $10^{-3}$ AdamW BCE
QCNN 50 128 $10^{-3}$ AdamW BCE
QLSTM 50 64 $10^{-3}$ AdamW BCE
QNN 50 512 $2\times 10^{-3}$ Adam BCE
QSVM Hinge
QASA 50 64 $10^{-3}$ AdamW BCE
LSTM 50 64 $10^{-3}$ AdamW BCE
Transformer 50 64 $10^{-3}$ AdamW BCE
[…]
Number of boosting iterations rather than epochs.
III-N Evaluation Framework
III-N1 Portfolio Construction
Model predictions are transformed to daily cross-sectional z-scores:
$\tilde{s}_{t}^{(i)}=\frac{s_{t}^{(i)}-\mu(s_{t})}{\sigma(s_{t})}$ (51)
Securities are filtered for trading feasibility (positive price/volume, no supervision flags). Long and short portfolios each contain the top/bottom $K=200$ ranked securities.
III-N2 Performance Metrics
The primary metric is the out-of-sample annualized Sharpe ratio:
$\text{SR}=\sqrt{252}\cdot\frac{\mu(r_{\text{LS}})}{\sigma(r_{\text{LS}})}$ (52)
where daily long-short returns are:
$r_{\text{LS},t}=\frac{1}{K}\sum_{i\in\mathcal{L}_{t}}r_{t+1}^{(i)}-\frac{1}{K}\sum_{j\in\mathcal{S}_{t}}r_{t+1}^{(j)}$ (53)
with $\mathcal{L}_{t}$ and $\mathcal{S}_{t}$ denoting long and short portfolios on day $t$.
III-N3 Statistical Significance
To assess performance reliability, we compute bootstrap confidence intervals for the Sharpe ratio following [ledoit2008robust]:
$\text{CI}_{95\%}(\text{SR})=[\hat{\text{SR}}-1.96\cdot\text{SE}(\text{SR}),\hat{\text{SR}}+1.96\cdot\text{SE}(\text{SR})]$ (54)
III-O Quantum Implementation Details
III-O1 Circuit Simulation
All quantum circuits are simulated using PennyLane [bergholm2018pennylane] with the default.qubit or lightning.qubit backends. Circuit depth is deliberately shallow ($L\leq 3$) to mitigate barren plateau phenomena [mcclean2018barren].
III-O2 Hardware-Efficient Ansatz
Entanglement structures follow nearest-neighbor connectivity compatible with near-term quantum devices:
$\text{CNOT}_{\text{ring}}=\prod_{j=1}^{n_{q}-1}\text{CNOT}_{j,j+1}\cdot\text{CNOT}_{n_{q},1}$ (55)
III-O3 Gradient Computation
PennyLane’s automatic differentiation integrates seamlessly with PyTorch [paszke2019pytorch], enabling end-to-end training of hybrid architectures through the parameter-shift rule (Equation 50).
III-P Reproducibility
III-P1 Random Seeds
For exact reproducibility, seeds should be fixed across NumPy, PyTorch, and PennyLane random number generators.
III-P2 Computational Resources
Quantum simulation scales exponentially with qubit count. With $n_{q}=8$ qubits, state vectors have dimension $2^{8}=256$, all neural network-based methods are trained 50 epochs, enabling efficient CPU-based simulation.
III-Q Conclusion
This methodology establishes a rigorous comparative framework for quantum-enhanced machine learning in quantitative finance. The unified evaluation protocol enables direct comparison across architecturally diverse approaches, providing a foundation for assessing the practical utility of quantum computing in financial prediction tasks.
III-R Computational Considerations for Quantum Simulation
A critical challenge in variational quantum algorithm research is the exponential scaling of classical simulation. For an $n_{q}$-qubit system, state vector simulation requires $O(2^{n_{q}})$ memory and $O(L\cdot 4^{n_{q}})$ time complexity per forward pass, where $L$ denotes circuit depth. Table II reports the empirical wall-clock time for a single forward-backward pass across all quantum architectures evaluated in this study, measured on Pannylane.
TABLE II: Computational Cost per Training Iteration (on RTX3090 GPU and 12th Gen Intel(R) Core(TM) i9-12900KS CPU)
Model Time/Iteration (s) Relative Cost
[…]
This computational overhead necessitates a carefully designed subsampling strategy to enable systematic comparison across architecturally diverse quantum models within practical time constraints.
IV Result
IV-A Experimental Design Rationale
IV-A1 Data Subsampling Strategy
Given the computational constraints of quantum circuit simulation (Section III-R), we adopt a stride-based subsampling scheme with $k=11$ to reduce the training set to approximately 9% of original trading days while preserving the full temporal span (2017–2021). This design choice balances two competing objectives:
  • Architectural coverage: Enabling fair comparison across 9 distinct model architectures, including 7 quantum variants with varying circuit depths and entanglement structures.
  • Statistical validity: Maintaining sufficient samples per trading day (approximately 400 stocks after Universe0 filtering) to ensure stable cross-sectional ranking signals.
The resulting dataset comprises approximately 100 training days ($\sim$40,000 stock-day samples) and 22 out-of-sample days ($\sim$8,800 stock-day samples), yielding a total quantum circuit evaluation count exceeding $10^{6}$ across all experiments.
IV-A2 Limitations and Scope
We acknowledge that the 22-day OOS period limits the statistical power for detecting small performance differences. The reported Sharpe ratios should be interpreted as relative performance indicators within this controlled benchmark setting, rather than definitive estimates of live trading performance. We defer extended backtesting on the full dataset to future work leveraging quantum hardware or GPU-accelerated simulation backends.
IV-B Out-of-Sample Performance Analysis
Table III presents the out-of-sample (OOS) Sharpe ratio for all evaluated models over a 22-day testing period. Among all approaches, the quantum models demonstrate a clear advantage in terms of risk-adjusted returns.
The proposed QTCNN achieves the highest OOS Sharpe ratio of $0.538\pm 0.042$, outperforming all baseline methods by a significant margin. This is followed by QNN ($0.467\pm 0.038$) and QLSTM ($0.333\pm 0.051$), both of which also surpass the classical Transformer baseline ($0.313\pm 0.036$). Notably, the top three performing models are all quantum-based architectures, suggesting that quantum feature encoding and variational circuits can effectively capture complex temporal patterns in financial time series.
In contrast, the classical LSTM yields the lowest Sharpe ratio of $0.044\pm 0.029$, indicating limited generalization capability in the OOS setting. The hybrid PQC+LGBM model ($0.122\pm 0.017$) also underperforms relative to end-to-end quantum models, implying that naive hybridization may not fully exploit the representational power of parameterized quantum circuits.
These results collectively validate our hypothesis that quantum convolutional architectures provide superior feature extraction for financial forecasting tasks, achieving approximately $72\%$ improvement over the best classical baseline (Transformer) in terms of Sharpe ratio.
Classical/Quantum Model Sharpe Ratio (OOS) OOS Days
Quantum QTCNN 0.538 $\pm$ 0.042 22
Quantum QNN 0.467 $\pm$ 0.038 22
Quantum QCNN 0.361 $\pm$ 0.298 22
Quantum QLSTM 0.333 $\pm$ 0.051 22
Classical Transformer 0.313 $\pm$ 0.036 22
Quantum QASA 0.266 $\pm$ 0.045 22
Quantum QSVM 0.131 $\pm$ 0.129 22
Quantum PQC+LGBM 0.122 $\pm$ 0.017 22
Classical LSTM 0.044 $\pm$ 0.029 22
TABLE III: Out-of-sample performance summary aggregated from project outputs.
IV-C Ablation Analysis
The performance hierarchy among quantum models provides insight into architectural design principles:
[…]
Quantum vs. Classical: The top-3 performing models are all quantum-based, with QTCNN achieving 72% improvement over the best classical baseline (Transformer). This suggests that quantum feature encoding provides complementary representational capacity for cross-sectional ranking tasks.
V Conclusion
V-A Summary
We proposed QTCNN, a hybrid quantum-classical architecture that combines temporal convolutional encoding with parameter-efficient quantum convolution for cross-sectional equity return prediction. The key innovation lies in the integration of a classical temporal CNN encoder that preserves sequential dynamics from multi-scale technical factors, followed by a hierarchical quantum convolutional structure with intra-layer parameter sharing. This design addresses two fundamental challenges in applying quantum machine learning to financial time-series: (1) the loss of temporal information inherent in dimensionality reduction via PCA, and (2) the trainability issues associated with over-parameterized variational quantum circuits.
Through comprehensive benchmarking on the JPX Tokyo Stock Exchange dataset, we systematically compared QTCNN against five alternative quantum architectures (QNN, QCNN, QLSTM, QASA, QSVM, PQC+LGBM) and three classical baselines (MLP, LSTM, Transformer). Our empirical results demonstrate that QTCNN achieves the highest out-of-sample Sharpe ratio of $0.538\pm 0.042$, representing a 72% improvement over the best classical baseline (Transformer: $0.313\pm 0.036$) and a 49% improvement over the standard QCNN ($0.361\pm 0.298$). Notably, the top three performing models are all quantum-based architectures, suggesting that quantum feature encoding provides complementary representational capacity for capturing nonlinear, high-dimensional relationships in financial data.
The ablation analysis further reveals that the performance advantage of QTCNN stems from two synergistic components: the temporal encoder that extracts semantically rich representations from sequential input, and the parameter-shared quantum convolution that enhances generalization while maintaining computational tractability. These findings provide actionable design principles for practitioners seeking to apply quantum machine learning to financial prediction tasks.
V-B Limitations and Future Work
Several limitations of this study warrant discussion and motivate directions for future research.
Computational constraints. The exponential scaling of quantum circuit simulation necessitated a stride-based subsampling strategy, reducing the dataset to approximately 9% of original trading days. Consequently, the out-of-sample evaluation period was limited to 22 trading days, which constrains the statistical power for detecting small performance differences and limits coverage of diverse market regimes. The reported Sharpe ratios should therefore be interpreted as relative performance indicators within this controlled benchmark setting, rather than definitive estimates of live trading performance. Future work will leverage GPU-accelerated simulation backends (e.g., lightning.gpu, cuQuantum) or actual quantum hardware access to enable full-scale backtesting across extended time horizons.
Market regime coverage. The subsampled dataset may not fully capture tail events, volatility clustering, or regime transitions that characterize real-world financial markets. Extended validation across multiple market cycles, including periods of high volatility and structural breaks, is necessary to establish the robustness of QTCNN under diverse market conditions.
Hardware noise characterization. All experiments in this study assume ideal, noiseless quantum simulation. Near-term quantum devices are subject to various noise sources including gate errors, decoherence, and measurement noise, which may significantly degrade model performance. Characterizing the noise resilience of QTCNN through simulation with realistic noise models (e.g., depolarizing channels, amplitude damping) and eventual deployment on quantum hardware remains an important direction for future investigation.
Feature engineering and alternative markets. The current study focuses on a predefined set of momentum, volatility, and volume-based features derived from price data. Incorporating alternative data sources such as sentiment indicators, macroeconomic variables, or order book information may further enhance predictive performance. Additionally, validating the generalizability of QTCNN across different equity markets (e.g., U.S., European, emerging markets) and asset classes (e.g., fixed income, commodities) would strengthen the practical applicability of the proposed framework.
Theoretical understanding. While our empirical results demonstrate the effectiveness of QTCNN, a rigorous theoretical analysis of its expressibility, trainability, and potential quantum advantage remains an open question. Future work will investigate the relationship between quantum circuit architecture and the inductive biases beneficial for financial prediction tasks.
In conclusion, this work establishes QTCNN as a promising quantum-classical hybrid architecture for cross-sectional equity prediction and provides a reproducible benchmark framework for evaluating quantum machine learning methods in quantitative finance. As quantum hardware continues to mature, we anticipate that the architectural principles identified in this study will inform the development of practical quantum-enhanced trading systems.
[Uncaptioned image] Chi-Sheng Chen received a bachelor degree in Electrophysics from National Chiao Tung University and a master’s degree in Biomedical Electronics and Bioinformatics from National Taiwan University. He is currently working at a financial AI startup and also serves as an AI project researcher at Harvard Medical School and Beth Israel Deaconess Medical Center. His research expertise includes time-series analysis, computer vision, brain-computer interfaces (BCI), electroencephalography (EEG), multimodal contrastive learning, and quantum machine learning. Chen has led the development of foundational models for EEG-based image generation, aiming to improve psychiatric treatment and the discovery of biomarkers for drug development. He has designed hybrid quantum-classical architectures for EEG encoding and pioneered quantum deep learning methods for contrastive learning between EEG and visual data. His research also focuses on medical speech processing and the integration of multimodal large language models (LLMs) for clinical and financial application.
[Uncaptioned image] Xinyu Zhang received the B.E. and the M.S. degree from the Information and Automatic College, Tianjin University of Science and Technology, Tianjin, China, in 2019 and 2022, separately, where he is currently pursuing the Ph.D. degree in intelligent system engineering with Luddy School of Informatics, Computing, and Engineering, Indiana University Bloomington, Bloomington, IN, USA. His main research includes quantum neural network, deep learning, advanced signal processing, and other disciplines such as information science and large language models (LLMs).
[Uncaptioned image] Rong Fu received the B.S. degree in Communication Engineering from Jiangsu Ocean University, Jiangsu, China, in 2018, and the M.S. degree in Electronic Information Engineering from the College of Information and Automation, Tianjin University of Science and Technology, Tianjin, China, in 2023. She is currently pursuing the Ph.D. degree in Intelligent Systems Engineering with the Luddy School of Informatics, Computing, and Engineering, Indiana University Bloomington, Bloomington, IN, USA. Her research interests include signal processing and deep learning.
[Uncaptioned image] Qiuzhe Xie received a bachelor’s degree in Electrical Engineering from Hefei University of Technology and a master’s degree in Electronics Engineering from National Cheng Kung University, followed by a Ph.D. in Electronics Engineering from National Taiwan University. His research expertise encompasses biosensors, advanced semiconductor processes, nanomaterials, and microfabrication techniques. He has spearheaded the development of nano-structural electrode arrays with ultra-thin dielectric stacking and leveraged quantum computing algorithms to enhance sensing precision. Furthermore, he has engineered CRISPR-based electrochemical sensors and developed innovative water purification systems for semiconductor manufacturing. Extending his work to practical application, he also focuses on the commercialization of medical devices, having validated business models for cardiosensor technologies through the NSF program.
[Uncaptioned image] Fan Zhang received the B.S. degree in Statistics from Beijing Normal University, Beijing, China, in 2016, the M.S. degree in statistics from National Cheng Kung University, Taiwan, in 2018, and the Ph.D. degree in statistics from Arizona State University, Tempe, AZ, USA, in 2024. She is currently an Assistant Professor with the Department of Mathematics, Boise State University, Boise, ID, USA. Her research interests include design and analysis of experiments, variable selection, uncertainty quantification, Bayesian modeling and inference, and quantum-enhanced machine learning.
[…]